Circular No. 7390

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

XTE J1118+480 M. Uemura and T. Kato, Kyoto University (KU); and H. Yamaoka,

Kyushu University, on behalf of the VSNET collaboration team, report the detection of the probable optical counterpart (unfiltered CCD mag 12.92) of XTE J1118+480 on Mar. 30.726 UT, using a KU 0.25-m telescope, at α = 11^{h}18^{m}10^{s}.79, δ = +48^{o}02′11′′.2 (equinox 2000.0; accuracy < 1′′; measured by Yamaoka with GSC-ACT stars), which is well within the error radius of the ASM position (IAUC 7389). Yamaoka further notes that a star with red mag 18.8 in the USNO A1.0 and A2.0 catalogues (A2.0 position end figures 10^{s}.851, 12′′.78) lies within 2′′ of this position, and no other object can be seen on images of the second Digital Sky Survey within 10′′. This suggests that this star is possibly the quiescent counterpart of XTE J1118+480. D. Buczynski, Condor Brow Observatory, confirmed the object at unfiltered CCD mag 12.91 on Mar. 30.863. Kato forwards magnitudes by K. Takamizawa, Saku-machi, Nagano (T-Max 400 film): fainter than 15.0–15.4 on 15 dates, 1994 Nov. 4–1999 Dec. 7.774; 2000 Jan. 8.743, 12.7; 27.595, 13.3; 28.604, 12.3; Feb. 27.617, 13.8; Mar. 24.481, 12.9. Visual magnitude estimate by O. Trondal, Oslo, Norway: Mar. 30.867, 12.5.

G. G. Pooley and E. M. Waldram, Mullard Radio Astronomy Observatory, University of Cambridge, write: “We obtained observations at 15 GHz with the Ryle Telescope of a field of diameter 15′, centered on the position of XTE J1118+480 (IAUC 7389), between Mar. 30.71 and 31.21 UT. Two sources were detected: NVSS J111820+475659 with flux density 5 mJy, and a source (not detected in NVSS) at α = 11^{h}18^{m}10^{s}.7, δ = +48^{o}02′14′′ (equinox 2000.0; uncertainty 3′′; rms noise 0.5 mJy) with flux density 6.2 mJy. The inferred variability or unusual spectrum of this source make it a candidate for the counterpart of the x-ray source.”

C. A. Wilson and M. L. McCollough, Marshall Space Flight Center, report for the Compton Gamma Ray Observatory BATSE Team: “The current outburst of XTE J1118+480 was first detected in 1-day CGRO/BATSE earth-occultation measurements at a level of 61 ± 16 mCrab (20–100 keV) on Mar. 26. We see evidence of a slow rise in flux since about Mar. 5. The January outburst (IAUC 7389) is of a much different character from the current outburst, having risen abruptly from < 20 mCrab (20–100 keV) on Jan. 4 to 78 ± 16 mCrab on Jan. 5 and reaching a peak of 110 ± 20 mCrab on Jan. 11; the outburst slowly declined and the source was undetectable by Jan. 29. The energy spectra from both the current outburst and the January outburst are well characterized by a power law with a photon index of 2.1 ± 0.1, and the source is visible up to 120 keV.”

2000 March 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green